# {textblock*}

### v2

@@ 2.1 Instanton geometries and actions
In case the underlying theory of quantum gravity allows for ETW branes, further instanton geometries exist. Let us briefly argue that this is a reasonable assumption: First, ETW branes can arise from shrinking extra dimensions [19]. Second, they appear naturally in type IIA string theory in the form of O8 planes. Finally, they are predicted to exist in any quantum gravity theory by the more general Cobordism Conjecture [20]. Recent studies of ETW brane examples include [21, 22, 23, 24, 25, 26, 17, 27, 28, 29, 30, 31]. [^1] We note in particular that, using an explicit regularization of the 4d ETW brane tension, actions can be obtained in a universal way from the 4d EFT[17, 31].
@@
We note that, locally, i.e. focusing on the vicinity of the equator of the sphere and of the waist of the de Sitter hyperboloid, this corresponds to the standard procedure of gluing the euclidean geometry to its analytic continuation. Globally, this is not the case since there is a boundary in the ‘past’ of the euclidean region and no corresponding deviation from a pure de Sitter space in the Lorentzian geometry. One could in principle improve on this by including a field, the profile of which signaling the proximity of the ETW brane, and analytically continuing such a more complete instanton solution. However, we believe that our approach is sufficiently precise under the assumption that any such field is stabilized away from the ETW brane, such that the ‘local’ analytic continuation described above can be trusted.
Thus, as just explained and as seen in the right column of Table 1, the nucleated on-shell space does not contain an ETW boundary. The ETW brane we used extends in the euclidean direction and marks the ‘beginning of time’, similarly to what is discussed in [34, 24]. Our spacelike ETW brane appears only in the off-shell region. To convince readers who doubt that spacelike ETW branes exist at all, we recall the situation in the more established Witten bubble-of-nothing [19] nucleation process: There, the locus where the $S^{1}$ shrinks effectively represents an ETW brane and part of it, in the of-shell region, must be spacelike for the decay to occur.
@@
and vanishes if the ETW brane tension approaches zero. This is intuitvely clear if one observes from (4) that, in the limit of vanishing tension, the radius of the ETW branes approaches $\ell_{dS}$, both branes approach the equator, and hence the instanton volume goes to zero.
@@ 2.2 Creation rates
Given the instanton geometry, one may compute the vacuum creation rates. For vacuum decays, the instanton action is related to the decay rate $\Gamma$ via [18]
@@ 3 Creation of toric universes
Such a creation from nothing process is, of course, also conceivable if the cosmological constant of the torus vanishes. But in this case the created universe would classically not expand and one only grow due to quantum fluctuations.
@@
While we belief that this puzzling result deserves further study, it naively appears that zero-tension ETW branes are able to nucleate (and also terminate) toric universes without any action cost. Torus universes would then undergo a continuous production and destruction cycle.
@@ Acknowledgments
We thank Iñaki Garcìa Etxebarria and Daniel Schiller for valuable discussions and comments. The work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).

### v3

@@ 2.1 Instanton geometries and actions
In case the underlying theory of quantum gravity allows for ETW branes, further instanton geometries exist. Let us briefly argue that this is a reasonable assumption: First, ETW branes can arise from shrinking extra dimensions [19]. Second, they appear naturally in type IIA string theory in the form of O8 planes. Finally, they are predicted to exist in any quantum gravity theory by the more general Cobordism Conjecture [20]. An early discussion of ETW branes with tension appears in [21]. Recent studies of ETW brane examples include [22, 23, 24, 25, 26, 27, 17, 28, 29, 30, 31, 32]. [^1] We note in particular that, using an explicit regularization of the 4d ETW brane tension, actions can be obtained in a universal way from the 4d EFT[17, 32].
@@
We note that, locally, i.e. focusing on the vicinity of the equator of the sphere and of the waist of the de Sitter hyperboloid, this corresponds to the standard procedure of gluing the euclidean geometry to its analytic continuation. Globally, this is not the case since there is a boundary in the ‘past’ of the euclidean region and no corresponding deviation from a pure de Sitter space in the Lorentzian geometry. One could in principle improve on this by including a field, the profile of which signals the proximity of the ETW brane, and analytically continuing such a more complete instanton solution. However, we believe that our approach is sufficiently precise under the assumption that any such field is stabilized away from the ETW brane, such that the ‘local’ analytic continuation described above can be trusted.
Thus, following the Coleman-de Luccia procedure and analytically continuing the instanton to Lorentzian signature, the nucleated on-shell space does not contain an ETW boundary, cf. the right column of Table 1. The ETW brane we used extends in the euclidean direction and marks the ‘beginning of time’, similarly to what is discussed in [35, 25]. Our spacelike ETW brane appears only in the off-shell region. To convince readers who doubt that spacelike ETW branes exist at all, we recall the situation in the more established Witten bubble-of-nothing [19] nucleation process: There, the locus where the $S^{1}$ shrinks effectively represents an ETW brane and part of it, in the of-shell region, must be spacelike for the decay to occur.
@@
and vanishes if the ETW brane tension approaches zero. This is intuitively clear if one observes from (4) that, in the limit of vanishing tension, the radius of the ETW branes approaches $\ell_{dS}$, both branes approach the equator, and hence the instanton volume goes to zero. We provide a path-integral saddle-point derivation of the purely geometric result (6) in Sect. 2.2.
@@
2.2 Explicit path integral analysis of the boundary process
Let us further study the boundary process by following [40, 41, 18, 3] and study the euclidean gravitational path integral. It has been proven to be a useful tool for computing vacuum decay rates as well as deriving thermodynamic properties of black holes.
Following e.g. [42, 43, 10], we consider a minisuperspace setup with metric
$ds^{2}=\frac{N^{2}}{q(t)}dt^{2}+q(t)d\Omega_{3}^{2}\,,$ (7)
a time interval $t\in[0,1]$ and a gauge where $N$ is constant. Setting $M_{P}=1$, the action reads
[…]
The first term is nothing but the Einstein-Hilbert action evaluated for the metric (7). Since our geometry has two boundaries, defined by $q(0)=q_{0}$ and $q(1)=q_{1}$, a standard GHY term is used in the evaluation. The second term in (8) is the contribution of an ETW brane with tension $T$ located at the $q=q_{0}$ boundary. Its form can be understood as follows: A domain wall with tension $T$ at a fixed physical position makes the following contribution to the action
$\int d^{4}x\sqrt{g}\,\delta\left(\frac{N}{\sqrt{q}}(t-t_{0})\right)T=\int d^{4}xNq\delta(t-t_{0})\left|\frac{\sqrt{q}}{N}\right|T=2\pi^{2}q(t_{0})^{3/2}\frac{N}{|N|}T\,.$ (9)
Setting $t_{0}=0$, the volume of the spherical brane eventually becomes $2\pi^{2}q_{0}^{3/2}$ and (8) follows.
We are interested in the amplitude
[…]
for an initial three-sphere of radius $q_{0}^{1/2}$ to evolve to a sphere with radius $q_{1}^{1/2}$. The wave function of the universe follows after integrating over all possible radii of the ETW brane:
[…]
We will evaluate it in the saddle point approximation. The saddle points of $G[q_{1},q_{0}]$ are determined by solutions of the equation of motion and the Hamiltonian constraint:
$\ddot{q}=-\frac{2}{\ell_{dS}^{2}}N^{2}\,,\qquad-\frac{3}{4N^{2}}\dot{q}^{2}+3=\frac{3q}{\ell_{dS}^{2}}+2T\delta(N)q_{0}^{3/2}\,.$ (12)
Using Dirichlet boundary conditions, as specified above, solutions of the former equation read
$\bar{q}(t)=-\frac{N^{2}t^{2}}{\ell_{dS}^{2}}+\left(\frac{N^{2}}{\ell_{dS}^{2}}+q_{1}-q_{0}\right)t+q_{0}\,.$ (13)
There are four values of $N$ ensuring that $\bar{q}$ is also a solution to the Hamiltonian constraint:
$N=\pm\ell_{dS}^{2}\left[\left(1-\frac{q_{0}}{\ell_{dS}^{2}}\right)^{1/2}\pm\left(1-\frac{q_{1}}{\ell_{dS}^{2}}\right)^{1/2}\right]\,,$ (14)
which for $q_{0},q_{1}<\ell_{dS}^{2}$ are real-valued. Following the intuition that a small sphere nucleates and grows, eventually going on-shell, we restrict attention to $q_{0}<q_{1}$. As discussed e.g. in [42, 43], in the case without ETW brane, the four solutions define saddle points potentially contributing to the path integral. There are various proposals for which saddle points are relevant. The two solutions with smaller $N^{2}$ describe geometries fitting within one hemisphere. By contrast, the solutions with larger $N^{2}$ corresponds to geometries containing the equator. Hartle and Hawking argue that the saddle point with smaller $N^{2}$ and $N>0$ defines the wavefunction, meanwhile the Linde/Vilenkin proposal corresponds to choosing the saddle point with the same $N^{2}$ but $N<0$. Following these arguments, the amplitude calculated in the saddle point approximation is given by
$G[q_{1};q_{0}]\propto\exp(\pm\left[4\pi^{2}\ell_{dS}^{2}\left(1-\frac{q_{0}}{\ell_{dS}^{2}}\right)^{3/2}-2\pi^{2}q_{0}^{3/2}T-4\pi^{2}\ell_{dS}^{2}\left(1-\frac{q_{1}}{\ell_{dS}^{2}}\right)^{3/2}\right])\,,$ (15)
with the positive and negative sign corresponding to the Hartle-Hawking and Linde/Vilenkin choice respectively. To find $\Psi$ from (11), we perform the integral over $q_{0}$, again in a saddle point approximation. Extremizing (15) with respect to $q_{0}$ yields the condition $T<0$ and
$q_{0}=\frac{4\ell_{dS}^{2}}{T^{2}\ell_{dS}^{2}+4}\,,$ (16)
which is precisely the result (4). Ultimately, we find for the wavefunction
$\Psi(q_{1})\propto\exp(\pm\left[4\pi^{2}M_{P}^{2}\ell_{dS}^{2}\sqrt{\frac{T^{2}\ell_{dS}^{2}}{T^{2}\ell_{dS}^{2}+4M_{P}^{4}}}-4\pi^{2}\ell_{dS}^{2}\left(1-\frac{q_{1}}{\ell_{dS}^{2}}\right)^{3/2}\right])\,.$ (17)
As explained in [42, 43], one obtains approximate solutions to the WDW equation in the regime $q_{1}>\ell_{dS}^{2}$ using the WKB matching procedure. Specifically, we analytically continue (17) to the on-shell regime and enforce reality in the Hartle-Hawking case or the outgoing-wave condition in the Linde-Vilenkin case:
$\Psi_{\rm HH}(q_{1})$ $\propto\exp(4\pi^{2}M_{P}^{2}\ell_{dS}^{2}\sqrt{\frac{T^{2}\ell_{dS}^{2}}{T^{2}\ell_{dS}^{2}+4M_{P}^{4}}}\,)\cos(4\pi^{2}\ell_{dS}^{2}\left(\frac{q_{1}}{\ell_{dS}^{2}}-1\right)^{3/2}-\frac{\pi}{4})\,,$ (18)
$\Psi_{\rm LV}(q_{1})$ $\propto\exp(-4\pi^{2}M_{P}^{2}\ell_{dS}^{2}\sqrt{\frac{T^{2}\ell_{dS}^{2}}{T^{2}\ell_{dS}^{2}+4M_{P}^{4}}}-i4\pi^{2}\ell_{dS}^{2}\left(\frac{q_{1}}{\ell_{dS}^{2}}-1\right)^{3/2})\,.$ (19)
For an analysis of the Swiss-Cheese instanton, one would need to leave the minisuperspace approximation, which is beyond the scope of this note.
@@ 2.3 Creation rates
Given the various instanton geometries from Sect. 2.1 or alternatively, the wavefunctions computed in Sect. 2.2, one may compute the vacuum creation rates. For vacuum decays, the instanton action is related to the decay rate $\Gamma$ via [18]
@@
Alternatively, one may derive the vacuum creation rates from the wavefunctions (18),(19). Following [42, 43], one finds that the creation rates agree with (23) and (24) for the Hartle-Hawking and Linde/Vilenkin wavefunction respectively.
@@ 3 Creation of toric universes
Such a creation from nothing process is, of course, also conceivable if the cosmological constant of the torus vanishes. But in this case the created universe would classically not expand and it can only grow due to quantum fluctuations.
@@
While we believe that this puzzling result deserves further study, it naively appears that zero-tension ETW branes are able to nucleate (and also terminate) toric universes without any action cost. Torus universes would then undergo a continuous production and destruction cycle.
@@ Acknowledgments
We thank Iñaki Garcìa Etxebarria and Daniel Schiller for valuable discussions and comments. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).